A man stands at the edge of a cliff and hits a golf ball into the ocean 101 m below. He hits the ball 38.5 m/s at an angle of 10 degrees above horizontal. Ignoring air resistance, how far does the ball travel horizontally before plunging into the ocean?
1
Expert's answer
2020-12-10T11:03:58-0500
"y=h-v\sin{10}t-0.5gt^2\\0=101-38.5\sin{10}t-0.5(9.8)t^2\\t=3.91\ s"

"x=vt\cos{10}=(38.5)(3.91)\cos{10}=148\ m"